# Characterization of the Cherenkov Photon Background for Low-Noise Silicon Detectors in Space

### v1

@@ Abstract
Future space observatories that seek to perform imaging and spectroscopy of faint astronomical sources will require ultra-low-noise detectors that are sensitive over a broad wavelength range. Silicon charge-coupled devices (CCDs), such as EMCCDs, skipper CCDs, multi-amplifier sensing (MAS) CCDs, and single-electron sensitive read out (SiSeRO) CCDs have demonstrated the ability to detect and measure single photons from X-ray energies to near the silicon band gap ( 1.1 m), making them candidate technologies for this application. In this context, we study a relatively unexplored source of low-energy background coming from Cherenkov radiation produced by energetic cosmic rays traversing a silicon detector. We present a model for Cherenkov photon production and absorption that is calibrated to laboratory data, and we use this model to characterize the residual background rate for ultra-low-noise silicon detectors in space. We study how the Cherenkov background rate depends on detector thickness, variations in solar activity, and the contribution of heavy cosmic ray species (). We find that for thick silicon detectors, such as those required to achieve high quantum efficiency at long wavelengths, the rate of cosmic-ray-induced Cherenkov photon production is comparable to other detector and astrophysical backgrounds. We apply our Cherenkov background model to simulated spectroscopic observations of extra-solar planets, and we find that thick detectors continue to outperform their thinner counterparts at longer wavelengths despite a larger Cherenkov background rate. Furthermore, we find that minimal masking of cosmic-ray tracks continues to maximize the signal-to-noise of very faint sources despite the existence of extended halos of Cherenkov photons.
@@ 1 INTRODUCTION
Silicon charge-coupled devices (CCDs) have emerged as the most popular detector technology for space-based astronomical observations ranging from soft X-rays to the very near-infrared (NIR; m). As sensitivity to faint astronomical sources continues to advance, intrinsic backgrounds in CCD detectors become an increasingly important consideration. For example, the search for potentially habitable Earth-like planets around Sun-like stars is one of the priority areas identified by the most recent Decadal Survey on Astronomy and Astrophysics (National Academies of Sciences, Engineering, and Medicine, 2021). NASA’s new flagship mission concept, the Habitable Worlds Observatory, seeks to perform direct imaging and spectroscopy of Earth-like planets around Sun-like stars (Clery, 2023). Such observations seek to detect “biosignatures”, aspects of a planet’s atmosphere or surface that may indicate the presence of life (Seager, 2013). However, direct spectroscopy of the atmospheres of Earth-like exoplanets presents an enormous technological challenge (Crill & Siegler, 2017; Crill, 2022). The light of the host star greatly outshines the reflected light of the planet, which may have an average rate of several photons per hour (Seager & Deming, 2010; Rauscher et al., 2019). In this regime of signal-starved observations and long integration times, ultra-low-noise detectors are vital. In particular, recent technology reports have stated the need for detectors that are sensitive at ultraviolet and visible wavelengths with a readout noise of e- rms/pix (Crill & Siegler, 2017; Crill, 2022). Given the precision required for these observations and the limited population of potentially observable exo-Earths (Mamajek & Stapelfeldt, 2024), it is critical to identify and characterize sources of detector, instrumental, environmental, and astrophysical backgrounds in order to accurately estimate the signal-to-noise ratio (SNR) of future exo-planet observations.
@@
Previous studies have considered several backgrounds that will affect the ability of a future space telescope to directly image and perform spectroscopy of exoplanets (Shaklan et al., 2013; Crill & Siegler, 2017; Rauscher et al., 2019; Crill, 2022; Steiger et al., 2024). These models often include intrinsic detector noise sources (e.g., readout noise, dark current, and clock-induced charge), as well as instrumental and astrophysical backgrounds (e.g., residual speckle light, zodiacal light, and exozodiacal light). Energetic cosmic rays are another familiar and well-characterized environmental background when operating in the intense radiation environment of space. The primary cosmic-ray track can be relatively easily identified through high charge occupancy, which allows for effective masking in science images (e.g., van Dokkum, 2001). While the cosmic-ray rate sets important limitations on instrument design and operation (e.g., radiation hardness, maximum exposure time, etc.), it is generally assumed to have a negligible impact on the rate of single-photon events that could be confused with faint astronomical signals. Here, we identify and characterize Cherenkov photon emission from energetic cosmic rays as a previously understudied source of background for photon-starved observations from space. Cherenkov photons that are emitted with energies near the silicon bandgap can travel a significant distance before depositing their energy in pixels that are outside conventional cosmic-ray masks. When combined across multiple exposures, Cherenkov photons represent a background that is indistinguishable from astronomical photons. In this paper, we demonstrate that for conventional cosmic-ray masking approaches, the Cherenkov photon background is comparable in magnitude to other astrophysical and detector backgrounds that are commonly considered.
@@
This paper is organized as follows. In Section 2, we briefly review the theory of Cherenkov radiation, while in Section 3 we describe our Cherenkov photon production model. In Section 4, we describe our simulations of the cosmic-ray environment and Cherenkov photon production at Earth–Sun Lagrange Point 2 (L2). In Section 5, we discuss the results of our simulations in the context of exoplanet sensitivity estimates, and we conclude in Section 6. This paper improves and extends previous work presented in Gaido et al. (2024).
@@ 2 CHERENKOV RADIATION
The Cherenkov effect (Čerenkov, 1934, 1937) is the process by which a charged particle traveling through a polarizable medium emits electromagnetic radiation (Figure 1). In order for Cherenkov emission to take place, the charged particle must be traveling faster than the phase velocity of light in the medium. In such cases, the electric dipoles of the molecules in the material are excited, emitting a constructively interfering conical electromagnetic wavefront analogous to a sonic boom (Jelley, 1955).
Frank & Tamm (1937) described the spectrum of electromagnetic radiation that is emitted as a function of particle charge, , and velocity relative to the speed of light in vacuum, , as well as the dielectric function of the medium, , which depends on the emitted frequency, . The differential number of photons produced per unit length per unit frequency is described as (Budini, 1953),
@@
The dielectric function depends on the characteristics of the material, and in silicon there is a sharp cutoff in Cherenkov photon energies at (Du et al., 2022). Depending on energy (i.e., the frequency, ), Cherenkov photons can be promptly absorbed or travel a considerable distance from the charged-particle track. After Cherenkov photons are emitted, the distance that they travel from the track of the charged particle before being absorbed is set by the photon attenuation length (Du et al., 2022),
Photons with energies have an absorption length in silicon that is smaller than the usual track width (m ), making them invisible in an image taken by a silicon detector such as a CCD. Furthermore, photons with energies 1.1 eV always escape the detector because their energy is less than the silicon bandgap and their absorption length is much larger than the dimensions of the detector. In the intermediate energy range between , a Cherenkov photon can travel an appreciable distance in silicon before being absorbed to generate a single electron-hole pair via the photoelectric effect, which we refer to as a “single-electron event” (Ramanathan & Kurinsky, 2020). The production of single-electron events from Cherenkov photons was recently identified as an important background in dark matter direct detection experiments using CCDs (Du et al., 2022). Cherenkov photon production in CCDs was recently measured in the laboratory (Giardino et al., 2022; Moroni et al., 2025) and considered as a background for space-based astronomical observations (Gaido et al., 2024).
@@ 3 CHERENKOV MODEL AND VALIDATION
To model Cherenkov emission in silicon, we first integrate Eq. 1 in frequency space by setting a constant value for the particle velocity, . The range of integration is determined by the photon frequencies that satisfy the Cherenkov emission condition for a given velocity (Eq. 1). Due to the nature of the dielectric function of silicon (Green, 2008; Du et al., 2022), integration always takes place in a single, continuous frequency interval,
@@
Our model for producing Cherenkov photons is applied to detailed simulations of cosmic-ray propagation and interaction in silicon detectors performed with the Geant4 simulation toolkit (Agostinelli et al., 2003). Geant4 has been used extensively to simulate the interactions of energetic particles with CCD detectors (e.g., Awaki et al., 2002; Ishiwatari et al., 2006; Aguilar-Arevalo et al., 2016, 2022; Arnquist et al., 2023). Using Geant4, we can simulate cosmic-ray interactions in silicon detectors with high spatial and physical fidelity. We generate cosmic-ray events by drawing from user-defined particle type, spatial, angular, and energy distributions. The primary particle is propagated through the detector material and each interaction is tracked. When secondary particles are generated they are are also tracked.
@@
We assume no external shielding around the detector in our Geant4 simulations. In reality, shielding by passive material in the instrument and spacecraft would reduce the energy of incoming particles, create secondary showers from interactions of the primary cosmic ray in the shielding material, and change the angular/energy distribution of events. Studies of shielding would require a detailed model of the instrument and spacecraft, and are thus beyond the scope of the current analysis.
@@
For each particle, we store each step of its propagation through the silicon detector. A propagation step contains the initial and final position, velocity, deposited energy, and particle type. The charge deposited at each step is diffused in the detector, simulating the drift of charge carriers generated in the bulk of the detector, and its spread is positively correlated with the depth of interaction (Barak et al., 2020). At each propagation step, we use Eq. 4 to calculate , randomly draw Cherenkov photons as described above, and propagate them by a distance (Eq. 3). Each photon either escapes the detector or is absorbed to produce a single-electron event. In this way, we self-consistently simulate images containing charged particle tracks and their corresponding low-energy halos of Cherenkov photons.
@@
We validate our model against cosmic-ray data collected with an ultra-low-noise silicon skipper CCD capable of counting single-electron events associated with Cherenkov emission. In particular, Giardino et al. (2022) and Moroni et al. (2025) present a laboratory data set of cosmic-ray muons observed with a 675-m-thick skipper CCD comprised of 6144866 15 m pixels. This device was designed at LBNL and fabricated at Teledyne DALSA Semiconductor on high-resistivity silicon. In order to study the low-energy halo around high-energy events, the skipper CCD was operated in “Smart Skipper” mode with both energy and region-of-interest readout (Chierchie et al., 2021). This readout configuration provides single-photon/single-electron resolution around energetic events while minimizing the readout time and dark current. Filtering for atmospheric muons, we can recreate these events in Geant4 by setting their incidence direction as measured in the original images and assuming that their energy is the characteristic atmospheric muon energy ( 4 GeV; Autran et al., 2018). By simulating several realizations of the same accepted muon events, we can compare the simulated and observed distributions of single-electron events as a function of the perpendicular distance to the muon track. We also measured the dark current event rate from the images and included this in our comparison with simulations. We find good agreement between data and simulations (Figure 2), which gives us confidence in extending our model to predict the Cherenkov photon background for silicon detectors in space.
@@ 4 SIMULATION OF SPACE ENVIRONMENT
We simulate the cosmic-ray environment at L2, the expected operational location for the Nancy Grace Roman Space Telescope (Spergel et al., 2015; Akeson et al., 2019) and the Habitable Worlds Observatory (Clery, 2023). The cosmic-ray environment at L2 closely follows the local interstellar cosmic-ray environment, which is composed primarily of protons with a steeply falling energy spectrum (Hillas, 2006; Particle Data Group, 2020). We obtain the cosmic-ray spectrum at L2 from the Cosmic-Ray Data Base (CRDB; Maurin et al., 2023), which provides the local interstellar spectrum (LIS) for fully ionized nuclei and includes a model for variation due to solar activity. Protons account for more than 90% of the total cosmic-ray flux at L2, while helium nuclei (alpha particles) are responsible for the majority of the remaining flux (Maurin et al., 2023). Heavier nuclei () are comparatively negligible in flux and contribute little to the Cherenkov photon rate (see Appendix A).
@@
Solar activity varies on an approximately 11-year cycle, modulating the cosmic-ray flux within the heliosphere. This modulation affects both the intensity and the energy distribution of cosmic rays observed at L2. To incorporate solar modulation, we follow the Force-Field approximation, as described in Zhu (2024). In this approximation, the effects of solar modulation can be parameterized by the solar modulation potential, , which ranges from GV (Maurin et al., 2023). This results in a higher particle rate at low solar activity and a lower rate at high solar activity, which is consistent with observations. Our model for the flux and composition of cosmic rays at L2 as a function of solar activity is shown in Figure 3. To quantify the effects of varying solar activity, we scale the number of simulated events per exposure according to the modulation shown in Figure 3. Solar activity shifts the minimum expected energy of cosmic rays as well; however, when considering only protons and alpha particles, the change in particle count dominates. As a reference, the minimum proton energy considered across all solar activity levels is 25 MeV.
@@
The silicon detectors are modeled as bulk silicon with a specified thickness and surface area. The interactions of cosmic rays with the silicon of the detector and the resulting Cherenkov photon production are simulated following the procedure described in Section 3. Primary cosmic rays are simulated uniformly on the surface of a sphere surrounding the detector yielding an isotropic angular distribution and a uniform spatial distribution. No surrounding material (e.g., instrument or spacecraft) is modeled. Minimal shielding (e.g., 10 mm of aluminum equivalent) will dominantly affect low-energy events and is only expected to change the high-energy proton flux by 20% at energies 25 MeV (e.g., Dawson et al., 2008). Considering this, the rates we present here are slightly higher than those measured for the James Webb Space Telescope operating at L2 (B. J. Rauscher & D. J. Fixsen, in prep.). We use Geant4 to generate a sufficient number of cosmic-ray events, such that each event is used only once per simulated set of exposures.

### v2

@@ Abstract
Future space observatories that seek to perform imaging and spectroscopy of faint astronomical sources will require ultra-low-noise detectors that are sensitive over a broad wavelength range. Silicon charge-coupled devices (CCDs), such as EMCCDs, skipper CCDs, multi-amplifier sensing (MAS) CCDs, and single-electron sensitive read out (SiSeRO) CCDs have demonstrated the ability to detect and measure single photons from X-ray energies to near the silicon band gap ( 1.1 m), making them candidate technologies for this application. In this context, we study a relatively unexplored source of low-energy background coming from Cherenkov radiation produced by energetic cosmic rays traversing a silicon detector. We present a model for Cherenkov photon production and absorption that is calibrated to laboratory data, and we use this model to characterize the residual background rate for ultra-low-noise silicon detectors in space. We study how the Cherenkov background rate depends on detector thickness, variations in solar activity, and the contribution of heavy cosmic ray species (). We find that for thick silicon detectors, such as those required to achieve high quantum efficiency at long wavelengths, the rate of cosmic-ray-induced Cherenkov photon production is comparable to other detector and astrophysical backgrounds. We apply our Cherenkov background model to simulated spectroscopic observations of extra-solar planets, and we find that thick detectors continue to outperform their thinner counterparts at longer wavelengths despite a larger Cherenkov background rate. Furthermore, we find that minimal masking of cosmic-ray tracks continues to maximize the signal-to-noise ratio of very faint sources despite the existence of extended halos of Cherenkov photons.
@@ 1 INTRODUCTION
Silicon charge-coupled devices (CCDs) have emerged as the most popular detector technology for space-based astronomical observations ranging from soft X-rays to the very near-infrared (NIR; m). As sensitivity to faint astronomical sources continues to advance, intrinsic backgrounds in CCD detectors become an increasingly important consideration. For example, the search for potentially habitable Earth-like planets around Sun-like stars is one of the priority areas identified by the most recent Decadal Survey on Astronomy and Astrophysics (Astro2020). NASA’s new flagship mission concept, the Habitable Worlds Observatory, seeks to perform direct imaging and spectroscopy of Earth-like planets around Sun-like stars (Clery:2023). Such observations seek to detect “biosignatures”, aspects of a planet’s atmosphere or surface that may indicate the presence of life (Seager:2013). However, direct spectroscopy of the atmospheres of Earth-like exoplanets presents an enormous technological challenge (Crill:2017; Crill:2022). The light of the host star greatly outshines the reflected light of the planet, which may have an average rate of several photons per hour (Seager:2010; Rauscher:2019). In this regime of signal-starved observations and long integration times, ultra-low-noise detectors are vital. In particular, recent technology reports have stated the need for detectors that are sensitive at ultraviolet and visible wavelengths with a readout noise of e- rms/pix (Crill:2017; Crill:2022). Given the precision required for these observations and the limited population of potentially observable exo-Earths (Mamajek:2024), it is critical to identify and characterize sources of detector, instrumental, environmental, and astrophysical backgrounds in order to accurately estimate the signal-to-noise ratio (SNR) of future exo-planet observations.
@@
Previous studies have considered several backgrounds that will affect the ability of a future space telescope to directly image and perform spectroscopy of exoplanets (Shaklan:2013; Crill:2017; Rauscher:2019; Crill:2022; Steiger:2024). These models often include intrinsic detector noise sources (e.g., readout noise, dark current, and clock-induced charge), as well as instrumental and astrophysical backgrounds (e.g., residual speckle light, zodiacal light, and exozodiacal light). Energetic cosmic rays are another familiar and well-characterized environmental background when operating in the intense radiation environment of space. The primary cosmic-ray track can be relatively easily identified through high charge occupancy, which allows for effective masking in science images (e.g., vanDokkum:2001). While the cosmic-ray rate sets important limitations on instrument design and operation (e.g., radiation hardness, maximum exposure time, etc.), it is generally assumed to have a negligible impact on the rate of single-photon events that could be confused with faint astronomical signals. Here, we identify and characterize Cherenkov photon emission from energetic cosmic rays as a previously understudied source of background for photon-starved observations from space. Cherenkov photons that are emitted with energies near the silicon bandgap can travel a significant distance before depositing their energy in pixels that are outside conventional cosmic-ray masks. When combined across multiple exposures, Cherenkov photons represent a background that is indistinguishable from astronomical photons. In this paper, we demonstrate that for conventional cosmic-ray masking approaches, the Cherenkov photon background is comparable in magnitude to other astrophysical and detector backgrounds that are commonly considered.
@@
This paper is organized as follows. In Section 2, we briefly review the theory of Cherenkov radiation, while in Section 3 we describe our Cherenkov photon production model. In Section 4, we describe our simulations of the cosmic-ray environment and Cherenkov photon production at Earth–Sun Lagrange Point 2 (L2). In Section 5, we discuss the results of our simulations in the context of exoplanet sensitivity estimates, and we conclude in Section 6. This paper improves and extends previous work presented in Gaido:2024.
@@ 2 CHERENKOV RADIATION
The Cherenkov effect (Cherenkov:1934; Cherenkov:1937) is the process by which a charged particle traveling through a polarizable medium emits electromagnetic radiation (Figure 1). In order for Cherenkov emission to take place, the charged particle must be traveling faster than the phase velocity of light in the medium. In such cases, the electric dipoles of the molecules in the material are excited, emitting a constructively interfering conical electromagnetic wavefront analogous to a sonic boom (Jelley:1955).
Frank:1937 described the spectrum of electromagnetic radiation that is emitted as a function of particle charge, , and velocity relative to the speed of light in vacuum, , as well as the dielectric function of the medium, , which depends on the emitted frequency, . The differential number of photons produced per unit length per unit frequency is described as (Budini:1953)
@@
The dielectric function depends on the characteristics of the material, and in silicon there is a sharp cutoff in Cherenkov photon energies at (Du:2022). Depending on energy (i.e., the frequency, ), Cherenkov photons can be promptly absorbed or travel a considerable distance from the charged-particle track. After Cherenkov photons are emitted, the distance that they travel from the track of the charged particle before being absorbed is set by the photon attenuation length (Du:2022)
Photons with energies have an absorption length in silicon that is smaller than the usual track width (m ), making them invisible in an image taken by a silicon detector such as a CCD. Furthermore, photons with energies 1.1 eV always escape the detector because their energy is less than the silicon bandgap and their absorption length is much larger than the dimensions of the detector. In the intermediate energy range between , a Cherenkov photon can travel an appreciable distance in silicon before being absorbed to generate a single electron-hole pair via the photoelectric effect, which we refer to as a “single-electron event” (Ramanathan_2020). The production of single-electron events from Cherenkov photons was recently identified as an important background in dark matter direct detection experiments using CCDs (Du:2022). Cherenkov photon production in CCDs was recently measured in the laboratory (Giardino:2022; moroni2025) and considered as a background for space-based astronomical observations (Gaido:2024).
@@ 3 CHERENKOV MODEL AND VALIDATION
To model Cherenkov emission in silicon, we first integrate Eq. 1 in frequency space by setting a constant value for the particle velocity, . The range of integration is determined by the photon frequencies that satisfy the Cherenkov emission condition for a given velocity (Eq. 1). Due to the nature of the dielectric function of silicon (GREEN20081305; Du:2022), integration always takes place in a single, continuous frequency interval,
@@
Our model for producing Cherenkov photons is applied to detailed simulations of cosmic-ray propagation and interaction in silicon detectors performed with the Geant4 simulation toolkit (Geant4:2002). Geant4 has been used extensively to simulate the interactions of energetic particles with CCD detectors (e.g., Awaki:2002; Ishiwatari:2006; DAMIC:2016; Oscura:2022; Arnquist:2023). Using Geant4, we can simulate cosmic-ray interactions in silicon detectors with high spatial and physical fidelity. We generate cosmic-ray events by drawing from user-defined particle type, spatial, angular, and energy distributions. The primary particle is propagated through the detector material and each interaction is tracked. When secondary particles are generated they are also tracked.
@@
We assume no external shielding around the detector in our Geant4 simulations. In reality, shielding by passive material in the instrument and spacecraft would reduce the energy of incoming particles, create secondary showers from interactions of the primary cosmic ray in the shielding material, and change the angular/energy distribution of events. Quantitative statements about the effects of shielding on the Cherenkov photon rate would require a detailed model of the instrument and spacecraft, and are thus beyond the scope of the current analysis.
@@
For each particle, we store each step of its propagation through the silicon detector. A propagation step contains the initial and final position, velocity, deposited energy, and particle type. The charge deposited at each step is diffused in the detector, simulating the drift of charge carriers generated in the bulk of the detector, and its spread is positively correlated with the depth of interaction (PhysRevLett.125.171802). At each propagation step, we use Eq. 4 to calculate , randomly draw Cherenkov photons as described above, and propagate them by a distance (Eq. 3). Each photon either escapes the detector or is absorbed to produce a single-electron event. In this way, we self-consistently simulate images containing charged particle tracks and their corresponding low-energy halos of Cherenkov photons.
@@
We validate our model against cosmic-ray data collected with an ultra-low-noise silicon skipper CCD capable of counting single-electron events associated with Cherenkov emission. In particular, Giardino:2022 and moroni2025 present a laboratory data set of cosmic-ray muons observed with a 675-m-thick skipper CCD comprised of 6144866 15 m pixels. This device was designed at LBNL and fabricated at Teledyne DALSA Semiconductor on high-resistivity silicon. In order to study the low-energy halo around high-energy events, the skipper CCD was operated in “Smart Skipper” mode with both energy and region-of-interest readout (Chierchie:2021). This readout configuration provides single-photon/single-electron resolution around energetic events while minimizing the readout time and dark current. Filtering for atmospheric muons, we can recreate these events in Geant4 by setting their incidence direction as measured in the original images and assuming that their energy is the characteristic atmospheric muon energy ( 4 GeV; Autran:2018). By simulating several realizations of the same accepted muon events, we can compare the simulated and observed distributions of single-electron events as a function of the perpendicular distance to the muon track. We also measured the dark current event rate from the images and included this in our comparison with simulations. We find good agreement between data and simulations (Figure 2), which gives us confidence in extending our model to predict the Cherenkov photon background for silicon detectors in space.
@@ 4 SIMULATION OF SPACE ENVIRONMENT
We simulate the cosmic-ray environment at L2, the expected operational location for the Nancy Grace Roman Space Telescope (Spergel:2015; Akeson:2019) and the Habitable Worlds Observatory (Clery:2023). The cosmic-ray environment at L2 closely follows the local interstellar cosmic-ray environment, which is composed primarily of protons with a steeply falling energy spectrum (hillas2006cosmic; 10.1093/ptep/ptaa104). We obtain the cosmic-ray spectrum at L2 from the Cosmic-Ray Data Base (CRDB; Maurin_2023), which provides the local interstellar spectrum (LIS) for fully ionized nuclei and includes a model for variation due to solar activity. Protons account for more than 90% of the total cosmic-ray flux at L2, while helium nuclei (alpha particles) are responsible for the majority of the remaining flux (Maurin_2023). Heavier nuclei () are comparatively negligible in flux and contribute little to the Cherenkov photon rate (see Appendix A).
@@
Solar activity varies on an approximately 11 yr cycle, modulating the cosmic-ray flux within the heliosphere. This modulation affects both the intensity and the energy distribution of cosmic rays observed at L2. To incorporate solar modulation, we follow the Force-Field approximation, as described in Zhu_2024. In this approximation, the effects of solar modulation can be parameterized by the solar modulation potential, , which ranges from GV (Maurin_2023). This results in a higher particle rate at low solar activity and a lower rate at high solar activity, which is consistent with observations. Our model for the flux and composition of cosmic rays at L2 as a function of solar activity is shown in Figure 3. To quantify the effects of varying solar activity, we scale the number of simulated events per exposure according to the modulation shown in Figure 3. Solar activity shifts the minimum expected energy of cosmic rays as well; however, when considering only protons and alpha particles, the change in particle count dominates. As a reference, the minimum proton energy considered across all solar activity levels is 25 MeV.
@@
The silicon detectors are modeled as bulk silicon with a specified thickness and surface area. The interactions of cosmic rays with the silicon of the detector and the resulting Cherenkov photon production are simulated following the procedure described in Section 3. Primary cosmic rays are simulated uniformly on the surface of a sphere surrounding the detector yielding an isotropic angular distribution and a uniform spatial distribution. No surrounding material (e.g., instrument or spacecraft) is modeled. Minimal shielding (e.g., 10 mm of aluminum equivalent) will dominantly affect low-energy events and is only expected to change the high-energy proton flux by 20% at energies 25 MeV (e.g., Dawson:2008). Considering this, the rates we present here are slightly higher than those measured for the James Webb Space Telescope operating at L2 (B. J. Rauscher & D. J. Fixsen, in prep.). We use Geant4 to generate a sufficient number of cosmic-ray events, such that each event is used only once per simulated set of exposures.